OpenStudy (firejay5):

In Items 5–7, f ind the inverse of each function. Use the definition to verify that the two functions are inverses. 5. f (x) = 6 - 3x 6. g (x) = x + 2 7. h(x) = -x + 5

OpenStudy (firejay5):

@Directrix @dtan5457

Directrix (directrix):

What does your book give as the definition to verify that the two functions are inverses.

Directrix (directrix):

I am still thinking about that altitude and temperature problem and what went wrong there.

Directrix (directrix):

@Firejay5 Seriously, I need the definition from your book.

OpenStudy (firejay5):

the definition is to change the function into the inverse: http://www.purplemath.com/modules/invrsfcn3.htm

OpenStudy (firejay5):

no it's to show you what I mean

Directrix (directrix):

5. f (x) = 6 - 3x y = 6 - 3x Interchange x and y x = 6 - 3y Okay, now solve that for y @Firejay5

OpenStudy (firejay5):

@Directrix I got y = x - 6/ -3

Directrix (directrix):

Use grouping symbols on your answer or your answer will be misconstrued.

Directrix (directrix):

I got y = -1/3 x + 2

OpenStudy (dtan5457):

Did you finish this? @Firejay5

OpenStudy (dtan5457):

My take on this is 1.y=-3x+6 inverse (which you already did is y=(x/-3)+2 to prove they are inverses you can do (f o g)x or (g o f)x f(x)=(x/-3)+2 g(x)=-3x+6 Let's say I do (f o g)x f(x)=(-3x+6/-3)+2 that becomes x-2+2 which equals x. if it equals x, it's the inverse.